OpenStudy (anonymous):

In the bet on a “single number” at roulette, there is chance 1/38 of winning. The bet pays 35 to 1; you can take this to mean that if you win the bet, your net gain is $35, and if you lose the bet then you lose $1 (that is, your net gain is -$1). You don’t need to know any more about roulette to solve this problem, but the description given above is what would happen if you were to bet $1 on a single number at a Las Vegas roulette table. Suppose you bet repeatedly on a single number; assume that the bets are independent of each other. Find the expected value of your net gain from 38 bets

OpenStudy (perl):

Expected value = (1/38) *35 + (37/38)(-1) = -1/19 which is around -$0.05 now multiply that by 38 so -1/19 * 38 = -2

OpenStudy (perl):

you will expect a net loss of 2 dollars on 38 bets